Verify The Following Identities:
1.cotα + tanα = cscα secα

2.cos θ cot θ
------------ - 1 = csc θ
1 - sin θ

Answer by stanbon(75887)

Verify The Following Identities:
1.cotα + tanα = cscα secα
----
cos/sin + sin/cos = (1/sin)(1/cos)
-----
(cos^2 + sin^2)/(sin*cos) = 1/(sin*cos)
============================================
2.cos θ cot θ
------------ - 1 = csc θ
1 - sin θ
-----------
(cos^2/sin)/(1-sin) -1 = 1/sin
-------
(1-sin^2)/[sin(1-sin)] = 1 + 1/sin
------
(1+sin)/sin = (sin+1)/sin
===============================
Cheers,
Stan H.
========================